OpenStudy (anonymous):

A projectile is fired horizontally from a 40m cliff at 24m/s . Find the speed and velocity of the projectile when it strikes the ground. find vx vy v

OpenStudy (anonymous):

if air resistance is negligable the: vx is just 24m/s vy is vy^2=uy^2+2as vy=sqrt(0+2*9.81*40) vy=28.01m/s v=sqrt(vx^2+vy^2) v=36.89m/s